OpenStudy (anonymous):

f(x) = x^2(3x^3 - 1) find the value of derivative of the function.

OpenStudy (australopithecus):

you can either expand it or use chain rule

OpenStudy (australopithecus):

I would probably just use chain rule

OpenStudy (australopithecus):

unless I was expected to simplify then I would just expand and use basic rules

OpenStudy (anonymous):

15x^4 -2x ?

OpenStudy (unklerhaukus):

\[f(x) = x^2(3x^3 - 1)\] \[f'(x) = (x^2)'(3x^3 - 1)+(x^2)(3x^3 - 1)'\]

OpenStudy (australopithecus):

f(x) = 3x^(6) - x^(2) f'(x) = 15x^(4) - 2x

OpenStudy (australopithecus):

or you can use chain rule like rhaukus is doing which is also very easy

OpenStudy (anonymous):

Thankx :))

OpenStudy (unklerhaukus):

\[f(x)=(x^2)(3x^3 - 1)=3x^5-x^2\] if you are going to expand the bracket make sure you get you indicies right

OpenStudy (unklerhaukus):

\[f'(x) = (x^2)'(3x^3 - 1)+(x^2)(3x^3 - 1)'\]\[\qquad=2x(3x^3-1)+x^2(9x^2)\]\[\qquad=6x^4-2x+9x^4\]